OpenStudy (anonymous):

can some 1 help me figure out this equation 2c-9=3c+9

OpenStudy (amistre64):

what do we have to do ?

OpenStudy (anonymous):

what letter are we working out?

OpenStudy (amistre64):

realize that the = sign means that whatever you do to one side has to be done to the other to keep them equal...

OpenStudy (amistre64):

2c-9=3c+9 ; lets -2c from both sides 2c-2c-9=3c-2c+9 -9 = c+9 ; now subtract 9 from both sides like this -9-9 = c+9-9 -18 = c+0 -18 = c

OpenStudy (anonymous):

thanks sooo much

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

x/3+2/3= -2x/3-2 this is confusing can i get sum help

OpenStudy (amistre64):

multiply everything by 3 to get rid of the fractions.... x+2 = -2x -6

OpenStudy (amistre64):

3x=-8 x = -8/3

OpenStudy (anonymous):

can someone explain how 2x+5=3x+1 has no solution

OpenStudy (anonymous):

am i just finding the like terms

OpenStudy (anonymous):

A boat travels 400 miles at a rate of 20 miles per hour. How many hours did it travel?